# Aligning Machine and Human Visual Representations across Abstraction Levels

### v3

@@ B.3.1 Principal Components Analysis of Representations
In Fig. 19 we show PCA projections of the space of image representations from four different models before and after fine-tuning on AligNet. It shows that the large-scale structure of the representational space becomes more structured and interpretable. Notice for example that animals (blue) and food (green) are clearly separated from artifacts such as furniture (red), musical instruments (purple) and clothing items (orange) which are much more clustered together. The same structuring effect — though to a lesser degree — can be seen for principal components three and four (see Fig. 20).
This increased structuring within the first few principal components is also reflected in the amount of variance explained by them. The third row of Fig. 19 clearly shows a notable increase of the proportion of the total variance of the representations that can be explained with only the first 5-15 principal components. It is also striking how different the global structure of the four models is before fine-tuning on AligNet, and how similar it becomes afterwards.
Figure 20: Projection of the representations of 85k ImageNet images from four different models (columns) onto their first two principal components both before (first row) and after fine-tuning on AligNet (second row). Similar to Fig. 19 but for components 3 & 4.
In Fig. 21 we further show projections of the representations using TriMap [Amid and Warmuth, 2020], which is a non-linear dimensionality reduction technique designed to better preserve both the global and the local structure of the data. While the coarse structure is similar to that of the first two PCs (see Fig. 19), the fine-grained structure reveals some interesting details. Note for example that in the TriMap projection the ViT-B exhibits a lot of fine-grained structure in the form of tiny clusters.
Figure 21: Projection of the representations of 85k ImageNet images from four different models (columns) onto 2 dimensions using TriMap [Amid and Warmuth, 2020] both before (first row) and after fine-tuning on AligNet (second row). Colors correspond to high-level categories of the images.
@@ B.3.2 Detailed results on alignment with the semantic hierarchy
In Fig. 22 we show the distribution of changes in relative distance between the representations of pairs of ImageNet images for four models. We measure this distance change relative to other changes (by $z$-scoring), because relative distances are more meaningful than absolute ones (e.g., scaling all representations by a factor of two would change absolute distances, but not relative ones), and absolute scales of all representations tend to increase during training.
In general, images from classes that come from the same superordinate category (such as two different species of bird) tend to end up relatively closer together, while those from different superordinate categories tend to end up farther apart. The exact distribution of changes depends on the prior representation structure of the models; the effect is stronger for models that had lower-quality initial representations.
In Fig. 23 we show a more detailed visualization, showing a matrix of average changes in the relative distances between a subset of the basic-level categories, grouped into several higher-level semantic categories (animals, clothes, food, furniture, musical instruments, sports equipment, and vehicles), which are plotted as blocks on the diagonals. The block structure of the changes is clear, with overall increases in similarity within these broad categories, and decreases in similarity between them.
@@
Figure 22: Changes in relative distances between stimuli reflect the superordinate category structure. Representations within the same superordinate category tend to end up relatively closer together than those in different superordinate categories.

### v4

@@ B.3.1 Principal Components Analysis of Representations
In Fig.˜22 we show PCA projections of the space of image representations from four different models before and after fine-tuning on AligNet. It shows that the large-scale structure of the representational space becomes more structured and interpretable. Notice for example that animals (blue) and food (green) are clearly separated from artifacts such as furniture (red), musical instruments (purple) and clothing items (orange) which are much more clustered together. The same structuring effect — though to a lesser degree — can be seen for principal components three and four (see Fig.˜23).
This increased structuring within the first few principal components is also reflected in the amount of variance explained by them. The third row of Fig.˜22 clearly shows a notable increase of the proportion of the total variance of the representations that can be explained with only the first 5-15 principal components. It is also striking how different the global structure of the four models is before fine-tuning on AligNet, and how similar it becomes afterwards.
Figure 23: Projection of the representations of 85k ImageNet images from four different models (columns) onto their first two principal components both before (first row) and after fine-tuning on AligNet (second row). Similar to Fig.˜22 but for components 3 & 4.
In Fig.˜24 we further show projections of the representations using TriMap [3], which is a non-linear dimensionality reduction technique designed to better preserve both the global and the local structure of the data. While the coarse structure is similar to that of the first two PCs (see Fig.˜22), the fine-grained structure reveals some interesting details. Note for example that in the TriMap projection the ViT-B exhibits a lot of fine-grained structure in the form of tiny clusters.
Figure 24: Projection of the representations of 85k ImageNet images from four different models (columns) onto 2 dimensions using TriMap [3] both before (first row) and after fine-tuning on AligNet (second row). Colors correspond to high-level categories of the images.
@@ B.3.2 Detailed results on alignment with the semantic hierarchy
In Fig.˜25 we show the distribution of changes in relative distance between the representations of pairs of ImageNet images for four models. We measure this distance change relative to other changes (by $z$-scoring), because relative distances are more meaningful than absolute ones (e.g., scaling all representations by a factor of two would change absolute distances, but not relative ones), and absolute scales of all representations tend to increase during training.
In general, images from classes that come from the same superordinate category (such as two different species of bird) tend to end up relatively closer together, while those from different superordinate categories tend to end up farther apart. Images that come from the same basic- or subordinate-level category move even closer together. The exact distribution of changes depends on the prior representation structure of the models; the effect is stronger for models that had lower-quality initial representations.
In Fig.˜26 we show a more detailed visualization, showing a matrix of average changes in the relative distances between a subset of the basic-level categories, grouped into several higher-level semantic categories (animals, clothes, food, furniture, musical instruments, sports equipment, and vehicles), which are plotted as blocks on the diagonals. The block structure of the changes is clear, with overall increases in similarity within these broad categories, and decreases in similarity between them.
@@
Figure 25: Changes in relative distances between stimuli reflect the superordinate category structure. Representations within the same superordinate category tend to end up relatively closer together than those in different superordinate categories. Representations of images in the same basic-level category end up closer together, and those in the same subordinate category move even closer together.
@@
Verifying the statistical significance of the changes. To quantify the statistical significance of the representational changes, we fit mixed-effects linear regressions accounting for the non-independence of the representational changes across different clusters. We use the same-subordinate condition as the reference level, as this offers the most stringent comparison for the same-basic condition (i.e. the effect size of the difference between the two conditions is visibly the smallest). All effects were generally highly significant, full results are presented in Tab.˜8.
                                                 Estimate Std. Error          df   t value  Pr(>|t|)(Intercept)                                      -2.25018    0.01612  2344.28109  -139.562   < 2e-16 ***relationshipSame basic category                   0.25678    0.04675  2344.28108     5.493  4.39e-08 ***relationshipSame superordinate category           1.62585    0.03372   619.98263    48.220   < 2e-16 ***relationshipDifferent superordinate categories    2.72148    0.03229   520.51180    84.275   < 2e-16 ***
(a) SigLIP-B
                                                 Estimate Std. Error         df  t value  Pr(>|t|)(Intercept)                                      -1.73190    0.01931 2280.89874  -89.667   < 2e-16 ***relationshipSame basic category                   0.22030    0.05600 2280.89872    3.934  8.62e-05 ***relationshipSame superordinate category           0.98145    0.04498  963.74930   21.818   < 2e-16 ***relationshipDifferent superordinate categories    2.13518    0.04384  870.21854   48.701   < 2e-16 ***---Signif. codes:  0 ‘***’ 0.001 ‘**’ 0.01 ‘*’ 0.05 ‘.’ 0.1 ‘ ’ 1
(b) ViT-B
                                                 Estimate Std. Error          df  t value  Pr(>|t|)(Intercept)                                      -2.18517    0.02733  1127.92111   -79.95   < 2e-16 ***relationshipSame basic category                   0.29318    0.07925  1127.92113     3.70  0.000226 ***relationshipSame superordinate category           1.75272    0.07325   819.54322    23.93   < 2e-16 ***relationshipDifferent superordinate categories    2.59777    0.07268   793.50251    35.74   < 2e-16 ***
(c) Dino ViT-B
                                                 Estimate Std. Error          df  t value  Pr(>|t|)(Intercept)                                      -1.11179    0.03202  1312.37727  -34.719   < 2e-16 ***relationshipSame basic category                   0.19989    0.09285  1312.37733    2.153    0.0315 *relationshipSame superordinate category           0.63283    0.08885  1098.25627    7.122  1.92e-12 ***relationshipDifferent superordinate categories    1.40199    0.08847  1079.24964   15.847   < 2e-16 ***
(d) Dino V2 ViT-B
Table 8: Statistical significance of the representation reorganization for the AligNet models, via mixed-effects linear regressions. (Note that same-subordinate category is the reference level; thus, the statistics on the intercept denote the magnitude and significance of the changes in relative distances for that category, and the other coefficients and statistics are relative to that change.
Reorganization across living and nonliving categories Given that the living vs. nonliving distinction is extremely salient to humans, and likely plays a large role in the odd-one-out-judgments, we also computed similar analyses to the above focusing specifically on this distinction. Specifically, we evaluated how representation distances change within- and between the living and nonliving categories. We show the results in Fig.˜27—the changes are qualitatively as expected. These differences are also all statistically significant for all models (all $t$s >2.7>2.7, all $p$s <0.01<0.01) in mixed-effects regressions.
Figure 27: Changes in relative distances between living and nonliving entities after alignment. Living entities tend to cluster more closely together after alignment, nonliving entities do so more weakly, and the two categories tend to move apart from one another.
Alignment across layers Since our alignment procedure focuses on aligning the final representations of the model, it is not obvious that we would see similar reorganization at earlier layers. We therefore performed similar representation-change analyses to the above across the representation layers of the model (using the mean representation across spatial positions). We show the results in Fig.˜28—generally, we see a growing degree of reorganization across the layers of the model.
Figure 28: Changes in relative distances between stimuli across the layers of the ViT-B/16 model. We see generally consistent directions of the effects across layers, with the magnitude of the changes increasing at higher layers. (Errorbars on points are bootstrap 95%-CIs. Curves are quadratic regression fits, to illustrate the overall trends.)
Comparing representation changes to UnaligNet Finally, we compare the representation changes in the UnaligNet ablation (see above) for ViT-B. We show the results in Fig.˜29. This reorganization is significantly different from that observed in the AligNet version (Tab.˜9). Surprisingly, the UnaligNet procedure results in greater divergence amongst images within the same basic- or subordinate-level category, compared to the original representations. To gain some qualitative insight into this effect, we compared pairs of images within the same subordinate category on which AligNet and UnaligNet maximally disagree (Fig.˜30). The results suggest that perhaps UnaligNet representations are focusing more on superficial features that distinguish images, such as color or style. We validate these results with a human experiment in the next section.
Figure 29: Changes in relative distances between stimuli in the UnaligNet ViT-B. We see reorganization at the basic and subordinate levels that is the opposite of what would be desired; viz. these categories are moving farther apart, with changes that are larger in magnitude than the changes in the superordinate categories.
                                                         Estimate   Std. Error         df t value  Pr(>|t|)(Intercept)                                               1.776e+00  2.470e-02  1.948e+03   71.89   < 2e-16 ***versionAligNet                                           -3.328e+00  2.013e-02  1.906e+04 -165.38   < 2e-16 ***relationshipSame basic                                   -3.288e-01  7.162e-02  1.948e+03   -4.59  4.71e-06 ***relationshipSame superordinate                           -2.606e+00  6.175e-02  1.076e+03  -42.20   < 2e-16 ***relationshipDifferent superordinate                      -1.877e+00  6.086e-02  1.015e+03  -30.84   < 2e-16 ***versionAligNet:relationshipSame basic                     4.447e-01  5.836e-02  1.906e+04    7.62  2.65e-14 ***versionAligNet:relationshipSame superordinate             3.232e+00  2.573e-02  1.906e+04  125.60   < 2e-16 ***versionAligNet:relationshipDifferent superordinate        3.846e+00  2.143e-02  1.906e+04  179.48   < 2e-16 ***
Table 9: Statistical significance of the differences in representation changes between UnaligNet and AligNet (interaction of changes by training version). All interactions are highly significant. (Note that the relationship is dummy coded, so the simple effect of versionAligNet shows the difference in the reorganization of the same-subordinate category between the model versions.)
(a) Golf cart
(b) Red panda
(c) Tape player
Figure 30: Pairs of same-category images for which the representations of AligNet and UnaligNet maximally disagree. AligNet groups these images more closely together due to their shared semantic content, whereas UnaligNet represents them fairly differently, perhaps on the basis of more superficial features like color vs. black & white, pixelization, style, etc.
B.3.3 Human validation of the representational differences between AligNet and UnaligNet
In order to further ascertain the origin of the differences between the AligNet and UnaligNet tuned versions of ViT-B, we compared representations within each ImageNet category to find the largest disagreements between the two models. Specifically, within each category we identified the pair of images for which the models’ representations most strongly disagreed in both directions: the pair that the AligNet model represented much more similary than UnaligNet, and the pair that UnaligNet represented much more similarly than AligNet (Note that in this context, since both models start from the same base model, it is equivalent to compare similarities directly or to compare the change in similarity from the base model.) We took the top 100 disagreements (in magnitude) in both directions across categories, for a total of 200 image pairs — half of which AligNet represents much more similarly than UnaligNet, and half of which UnaligNet represents much more similarly than AligNet.
We then ran a human online experiment via Prolific ($N=49$ participants) testing which images participants rated (on a 1-5 Likert scale) as semantically more similar in content, when explicitly instructed to ignore superficial visual features. We find strong alignment between the participants’ judgments and the AligNet model — the participants consistently gave higher similarity ratings to the images that AligNet represented as more similar. This effect was highly significant (paired $t$-test: $t(48)=20.260,p<0.001$); in fact, every participant gave numerically higher similarity scores to the pairs AligNet rated as more similar. Most importantly, the correlation between model- and human-judged similarity is strongly positive for the aligned model (Spearman =0.641, p<.001) and negative for the unaligned model (Spearman =–0.122, p = 0.087).
These strong results support our claim that AligNet is uniquely changing the model’s representation structure—even within low-level categories—to align more closely with human semantic knowledge.
Figure 31: Human ratings of semantic similarity on image pairs on which AligNet and UnaligNet maximally disagree. The human participants consistently rate the pairs which AligNet represents more similarly as in fact more semantically similar. Shown on the right are the three image pairs that received the lowest human similarity ratings, all of which contradict UnaligNet’s prediction of high similarity.